OpenStudy (lgbasallote):

\[\left[\begin{array}{cc} 0 & 1 \\ 2 & 3 \end{array} \right] \times \left[\begin{array}{cc} 3 & 2\\ 1 & 0 \end{array} \right]?\]

OpenStudy (lgbasallote):

let me rephrase my question...

mathslover (mathslover):

wait i am typing the answer , it will take time

OpenStudy (lgbasallote):

\[\left[\begin{array}{cc} a & b \\ c & d \end{array} \right] \times \left[\begin{array}{cc} e & f\\ g& h \end{array} \right]?\]

mathslover (mathslover):

yes that is what i am going to type

OpenStudy (lgbasallote):

haha lol great minds :D

OpenStudy (lgbasallote):

i just figured if someone is gonna direct answer me...if i make it into variables i'll get what happened

OpenStudy (lgbasallote):

but i still do appreciate step-by-step...

OpenStudy (anonymous):

ae+bg af+bh ce+dg cf+dh ?

OpenStudy (anonymous):

2 by 2 matrices

mathslover (mathslover):